### v1

# Nearly Optimal Sample Complexity of Offline KL-Regularized Contextual Bandits under Single-Policy Concentrability

Abstract
KL-regularized policy optimization has become a workhorse in learning-based decision making, while its theoretical understanding is still very limited. Although recent progress has been made towards settling the sample complexity of KL-regularized contextual bandits, existing sample complexity bounds are either $\widetilde{O}(\epsilon^{-2})$ under single-policy concentrability or $\widetilde{O}(\epsilon^{-1})$ under all-policy concentrability. In this paper, we propose the first algorithm with $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability for offline contextual bandits. Our algorithm is designed for general function approximation and based on the principle of pessimism in the face of uncertainty. The core of our proof leverages the strong convexity of the KL regularization, and the conditional non-negativity of the gap between the true reward and its pessimistic estimator to refine a mean-value-type risk upper bound to its extreme. This in turn leads to a novel covariance-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class. The near-optimality of our algorithm is demonstrated by an $\widetilde{\Omega}(\epsilon^{-1})$ lower bound. Furthermore, we extend our algorithm to contextual dueling bandits and achieve a similar nearly optimal sample complexity.
1 Introduction
Due to the data-hungry and instable nature of reinforcement learning (RL), divergences that are straightforward to estimate via Monte Carlo methods or amenable to constrained optimization stand out from numerous candidates (Rényi, 1961; Müller, 1997; Basseville, 2013) as regularizers; the former family of which is typically $f$-divergence because any of them is an expectation under one policy (thus easy for on-policy estimation) (Levine, 2018; Levine et al., 2020), and the latter are naturally cases of Bregman divergence (Bregman, 1967) induced by strongly convex functions because they are suitable for scalable optimization algorithms over the probability simplex by design (Lattimore and Szepesvári, 2020; Lattimore, 2024). In particular, Kullback-Leibler (KL) divergence is the only one at the intersection of Bregman divergence and $f$-divergence (Jiao et al., 2014, Theorem 5), indicating that it is uniquely favorable among common choices from both computational and statistical aspects. Therefore, the KL-regularized RL objective is arguably the most popular one in practice:
$J(\pi)=\mathbb{E}_{\pi}[r]-\frac{1}{\eta}\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}}),$ (1.1)
where $r$ is a reward function, $\pi$ is the policy to be optimized, $\pi^{\mathsf{ref}}$ is a reference policy, $\mathsf{KL}({\pi}\|{\pi^{\mathsf{ref}}})$ is the reverse KL divergence, and $\eta>0$ is the inverse temperature. When $\pi^{\mathsf{ref}}$ is chosen to be the uniform distribution, the objective function reduces to the (Shannon) entropy-regularized RL objective that encourages diverse actions and enhances robustness (Williams, 1992; Ziebart et al., 2008; Ziebart, 2010; Levine and Koltun, 2013; Fox et al., 2015; Levine et al., 2016; Haarnoja et al., 2018; Richemond et al., 2024; Liu et al., 2024). The KL-regularized objective has also been widely used in the RL fine-tuning of large language models (LLMs) (Ouyang et al., 2022; Rafailov et al., 2023, 2024), where $\pi^{\mathsf{ref}}$ is a pretrained LLM.
There has also been a surge of interest in understanding the principle behind KL-regularized RL. Ahmed et al. (2019); Liu et al. (2019) studied by ablation the effect of entropy regularization on the stability and policy improvement in policy optimization, the regret of which has been rigorously settled under the classic online mirror descent framework (Cai et al., 2020; He et al., 2022; Ji et al., 2023). Neu et al. (2017) unified popular KL-regularized policy optimization algorithms under a convex optimization framework, but the interplay of which with data were left untouched. A series of work (Geist et al., 2019; Vieillard et al., 2020; Kozuno et al., 2022) then analyzed the sample complexity of algorithms using KL w.r.t. the previous iteration or/and entropy regularizers with improved dependence on the effective horizon in discounted Markov decision processes (MDPs). However, the metric of suboptimality in all these studies is still with respect to the unregularized reward maximization objective, and its optimal sample complexity is $O(\epsilon^{-2})$ for finding the $\epsilon$-optimal policy.
Several recent studies (Xiong et al., 2024; Xie et al., 2024; Zhao et al., 2024) switched the focus to suboptimality guarantee with respect to the regularized objective in (1.1). In particular, Xiong et al. (2024) proposed an Offline GSHF algorithm via the principle of pessimism in the face of uncertainty, and proved $O(\epsilon^{-2})$ sample complexity under single-policy concentrability (See Section 2.1 for detailed definitions of concentrability.). On the other hand, the sharp analysis in Zhao et al. (2024) yields the optimal sample complexity $O(\epsilon^{-1})$, but requires all-policy concentrability (Zhao et al., 2024, Definition 2.6), i.e., the data-generating policy $\pi^{\mathsf{ref}}$ is required to cover the entire function class for all possible policies. Xie et al. (2024) studied token-level Markov decision processes (MDPs) and proposed a KL-regularized RL algorithm named XPO, which achieves $O(\epsilon^{-2})$ sample complexity under all-policy concentrability. In fact, even for offline contextual bandits (Foster and Rakhlin, 2023, Chapter 3) with $\mathrm{i.i.d.}$ data, which is among the simplest setting for offline decision making, existing sample complexity under KL-regularization are still primitive. Thus, a natural question arises:
Is the $\widetilde{O}(\epsilon^{-1})$ sample complexity w.r.t. 1.1 achievable for offline KL-regularized contextual bandits under single-policy concentrability?
We answer this question affirmatively by proposing a novel yet minimalist approach based on the principle of pessimism (Jin et al., 2021). Our algorithm can deal with general reward function which belongs to a reward function class with finite covering entropy. The optimal sample complexity is also due to our proof technique, which leverages the strong convexity of KL regularization and the conditional non-negativity of the gap between the true reward and its pessimistic estimator to refine a mean-value-type risk upper bound to its extreme. This in turn leads to a novel covariance-based analysis, effectively bypassing the need for uniform control over the discrepancy between any two functions in the function class.
Table 1: Comparison of sample complexity for finding $\epsilon$-optimal policy for offline contextual bandits. $\mathrm{polylog}(\epsilon^{-1})$ and the dependence of other parameters are omitted here. C(D)Bs stands for contextual (dueling) bandits. The last column indicates whether the bounds hold under single-policy concentrability.
Reference [^1] The four existing upper bounds are adapted from the implicit form in Xiong et al. (2024, Theorem 3.1) and Zhao et al. (2024, Theorem 3.3 and Theorem 4.4.). See Section 2.2 for details. CBs CDBs Single?
Xiong et al. (2024) $\widetilde{O}(\epsilon^{-2})$ $\widetilde{O}(\epsilon^{-2})$
Zhao et al. (2024) $\widetilde{O}(\epsilon^{-1})$ $\widetilde{O}(\epsilon^{-1})$
This work $\widetilde{O}(\epsilon^{-1})$ $\widetilde{O}(\epsilon^{-1})$
Lower bounds $\widetilde{\Omega}(\epsilon^{-1})$ $\widetilde{\Omega}(\epsilon^{-1})$  [^2] The lower bounds in terms of $\epsilon$ are the same under both single- and all-policy concentrability, which are detailed in Remarks 2.11 and 3.8 respectively.
1.1 Contributions
  • We propose a new algorithm dubbed KL-PCB and develop a novel proof technique that simultaneously utilizes the curvature of KL-regularized objectives and integrates pessimism with a novel covariance-based argument, for deriving its $\widetilde{O}(\epsilon^{-1})$ sample complexity under single-policy concentrability.
  • To demonstrate the versatility of our techniques beyond absolute reward feedback, we also extend our results to the offline contextual dueling bandit (CDB) setting, and provide a tight sample complexity analysis.
  • The matching lower bounds (up to $\mathrm{polylog}$ factors) in both settings suggest that our sample complexity results are nearly optimal, which renders a complete picture of the sample complexity for KL-regularized offline contextual under single policy concentrability.
For the ease of comparison, we summarize the relevant sample complexity results along with the corresponding coverage assumptions in Table 1.
1.2 Additional Related Work
We review two additional lines of theoretical progress that are closely related to our algorithm design and analysis.
Pessimism in Offline RL
The principle of pessimism has been underpinning offline RL for both the tabular (Rashidinejad et al., 2021) and function approximation (Jin et al., 2021) settings under the name of lower confidence bound (LCB). For contextual bandits, it is behind the adaptively optimal sample complexity analysis (Li et al., 2022). Shi et al. (2022) proposed a LCB-based model-free algorithm for tabular RL with near-optimal guarantee. Jin et al. (2021); Xiong et al. (2022); Di et al. (2023) utilized LCB in conjunction with the classic least-square value iteration paradigm to derive $\widetilde{O}(\epsilon^{-2})$ sample complexity results for model-free RL with function approximation. The line of work from Rashidinejad et al. (2021); Xie et al. (2021b) to Li et al. (2024) settled the sample complexity of tabular model-based RL via pessimistic estimators exploiting the variance information. It is also possible to leverage the idea of pessimism to design model-based algorithms under general function approximation that are at least statistically efficient (Xie et al., 2021a; Uehara and Sun, 2021; Wang et al., 2024).
However, in terms of risk decomposition, to the best of our knowledge, none of these pessimism-based analyses really goes beyond the performance difference lemma (Foster and Rakhlin, 2023, Lemma 13) or simulation lemma (Foster and Rakhlin, 2023, Lemma 23); both of which are not able to capture the strong concavity of KL-regularized objectives even in the bandit setting. The algorithmic idea of using pessimistic least-square estimators under general function approximation in Jin et al. (2021); Di et al. (2023) is similar to ours, but their suboptimality gap is bounded by the sum of bonuses, the direct adaptation of which to our objective cannot lead to the desired sample complexity.
Offline Contextual Dueling Bandits
CDBs (Dudík et al., 2015) is the contextual extension of dueling bandits in classic literature of online learning from pairwise comparisons (Yue et al., 2012; Zoghi et al., 2014). Since the empirical breakthrough of preference-based RL fine-tuning of LLMs (Ouyang et al., 2022), the theory of offline CDBs has received more attention under linear (Zhu et al., 2023; Xiong et al., 2024) and general (Zhan et al., 2022; Zhao et al., 2024; Song et al., 2024; Huang et al., 2024) function approximation. Preference models without stochastic transitivity (Munos et al., 2023; Ye et al., 2024; Wu et al., 2024; Zhang et al., 2024) are beyond the scope of this work, viz., our preference labels are assumed to follow the Bradley-Terry Model (Bradley and Terry, 1952).
Notation
Boldfaced lower case (resp. upper case) letters are reserved for vectors (resp. matrices). The sets ${\mathcal{S}}$ and $\mathcal{A}$ are assumed to be countable throughout this paper to avoid measure-theoretic arguments. For nonnegative sequences $\{x_{n}\}$ and $\{y_{n}\}$, we write $x_{n}=O(y_{n})$ if $\limsup_{n\to\infty}{x_{n}}/{y_{n}}<\infty$, and write $y_{n}=\Omega(x_{n})$ if $x_{n}=O(y_{n})$. We further employ $\widetilde{O}(\cdot)$ and $\widetilde{\Omega}(\cdot)$ to hide $\mathrm{polylog}$ factors. Given a positive definite $\bm{\Sigma}\in\mathbb{R}^{d\times d}$ and $\mathbf{x}\in\mathbb{R}^{d}$, we denote the vector’s Euclidean norm by $\|\mathbf{x}\|_{2}$ and define $\|\mathbf{x}\|_{\bm{\Sigma}}=\sqrt{\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}}$. For coutable $\mathcal{X}$ and $\mathcal{Y}$, we denote the family of probability kernels from $\mathcal{X}$ to $\mathcal{Y}$ by $\Delta(\mathcal{Y}|\mathcal{X})$. For a pair of probability measures $P\ll Q$ on the same space, the KL divergence from $P$ to $Q$ is ${\mathsf{KL}}\left(P\|Q\right)\coloneqq\int\log({{\,\mathrm{d}}P}/{{\,\mathrm{d}}Q}){\,\mathrm{d}}P$, and their total variation (TV) distance is ${\mathsf{TV}}\left(P\|Q\right)\coloneqq 0.5\int|{\,\mathrm{d}}P-{\,\mathrm{d}}Q|$. For $f:\mathcal{X}\to\mathbb{R}$, its infinity norm is denoted by $\|f\|_{\infty}\coloneqq\sup_{x\in\mathcal{X}}|f(x)|$.
2 KL-Regularized Contextual Bandits
In this section, we propose a pessimism-based algorithm, dubbed PCB-KL, for offline KL-regularized contextual bandits. In the following subsections, we also showcase our key technical novelty which couples the pessimism of our reward estimator to the non-trivial curvature property of KL regularization.
2.1 Problem Setup
We consider contextual bandit, which is denoted by a tuple $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$. Specifically, ${\mathcal{S}}$ is the context space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. In the offline setting, the agent only has access to an $\mathrm{i.i.d.}$ dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$. Here $s_{i}\in{\mathcal{S}}$ is the state sampled from a distribution $\rho$, $a_{i}\in\mathcal{A}$ is the action taken from a behavior policy, and $r_{i}$ is the observed reward given by $r_{i}=r(s_{i},a_{i})+\varepsilon_{i}$, where $\varepsilon_{t}$ is $1$-sub-Gaussian (Lattimore and Szepesvári, 2020, Definition 5.2). In this work, we consider the KL-regularized objective
$J(\pi)\coloneqq\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}r(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$ (2.1)
where $\pi^{\mathsf{ref}}$ is a known reference policy and the “inverse temperature” $\eta$ controls the intensity of regularization. For simplicity, we assume that $\pi^{\mathsf{ref}}$ is also the behavior policy that generates the dataset $\mathcal{D}$. This type of “behavior regularization” has been also been studied in Zhan et al. (2022). The unique optimal policy $\pi^{*}\coloneqq\mathop{\mathrm{argmax}}_{\pi\in\Delta(\mathcal{A}|{\mathcal{S}})}J(\pi)$ is in the following closed-form (See, e.g., Zhang 2023, Proposition 7.16)
$\pi^{*}(\cdot|s)\propto\exp\big{(}\eta\cdot r(s,\cdot)\big{)},\forall s\in{\mathcal{S}}.$ (2.2)
A policy $\pi$ is said to be $\epsilon$-optimal if $J(\pi^{*})-J(\pi)\leq\epsilon$ and the goal of the agent is to find one such policy based on $\mathcal{D}$. To ensure $\epsilon$-optimality is achievable, we assume that $r$ lies in a known function class $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to[0,1])$, from which the agent obtains her estimator $\widehat{f}$.
Assumption 2.1 (General Function Approximation).
There exists a known function class $\mathcal{F}:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ such that $\exists\ f^{*}\in\mathcal{F}$ such that $f^{*}=r$.
We also need a standard condition to control the complexity of $\mathcal{F}$ through the notion of covering number (Wainwright, 2019, Definition 5.1).
Definition 2.2 ($\epsilon$-net and covering number).
Given a function class $\mathcal{G}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$, a finite set $\mathcal{G}(\epsilon)\subset\mathcal{G}$ is an $\epsilon$-net of $\mathcal{G}$ w.r.t. $\|\cdot\|_{\infty}$, if for any $g\in\mathcal{G}$, there exists $g^{\prime}\in\mathcal{G}(\epsilon)$ such that $\|g-g^{\prime}\|_{\infty}\leq\epsilon$. The $\epsilon$-covering number is the smallest cardinality $\mathcal{N}_{\mathcal{G}}(\epsilon)$ of such $\mathcal{G}(\epsilon)$.
Assumption 2.3.
For any $\epsilon_{c}>0$, the $\epsilon_{c}$-covering number $\mathcal{N}_{\mathcal{F}}(\epsilon_{c})$ of $\mathcal{F}$ is $\mathrm{poly}(\epsilon_{c}^{-1})$.
Remark 2.4.
When $\mathcal{F}$ is the class of linear functions of dimension $d$ and radius $R$, the covering number of $\mathcal{F}$ is given by $\mathcal{N}_{\mathcal{F}}=O((1+R\epsilon^{-1})^{d})$ (Jin et al., 2020, Lemma D.6), indicating that the Assumption 2.3 on $\mathcal{N}_{\mathcal{F}}$ is mild.
Concentrability
The data quality of $\mathcal{D}$ collected by $\pi^{\mathsf{ref}}$ is typically characterized by concentrability in offline RL (Farahmand et al., 2010; Chen and Jiang, 2019; Jiang and Xie, 2024). In the traditional definition based on density ratio (See, e.g., Zhan et al. 2022, Table 1), a strong concentrability condition intuitively corresponds to a large $\text{supp}(\pi^{\mathsf{ref}})$, which reduces the agent’s difficulty of learning since (with high probability) $\widehat{\pi}$ can hardly underfit on many state-action pairs. The two extreme cases of concentrability coefficients, all- and single-policy concentrability, intuitively correspond to (1) $\text{supp}(\pi^{\mathsf{ref}})$ covers all possible inputs; and (2) $\text{supp}(\pi^{\mathsf{ref}})$ only subsumes $\text{supp}(\pi^{*})$. Our definition of concentrability, however, is different from the density ratio-based one, because we take the function class $\mathcal{F}$ into account. In detail, we first introduce $D^{2}$-divergence as follows.
Definition 2.5.
Given a function class $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and a fixed policy $\pi$, define the $D^{2}$-divergence $D^{2}_{\mathcal{F}}((s,a);\pi)$ as
$\sup_{f,g\in\mathcal{F}}\frac{\big{(}f(s,a)-g(s,a)\big{)}^{2}}{\mathbb{E}_{(s^{\prime},a^{\prime})\sim\rho\times\pi}[(f(s^{\prime},a^{\prime})-g(s^{\prime},a^{\prime}))^{2}]}.$
The “Eluder dimension”-type Definition 2.5 is directly inspired by Di et al. (2023); Zhao et al. (2024), the intuition behind which is that given $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, a small $D^{2}$-divergence indicates that for two functions $f$ and $g$, if they are close under the behavior policy $\pi$, then they will also be close on such pair $(s,a)$. Therefore, the $D^{2}$-divergence quantifies how well the estimation on dataset collected by the behavior policy $\pi$ can be generalized to a specific state-action pair. We are now ready to define the two notions of concentrability conditions.
Assumption 2.6 (All-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D>0$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})$.
Assumption 2.6 indicates that the errors on any state-action pairs can be bounded by the error on the samples from $\rho\times\pi$ up to a factor $D$.
Assumption 2.7 (Single-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D_{\pi^{*}}>0$ such that $D_{\pi^{*}}^{2}=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})]$.
Single-policy concentrability indicates that the errors on the distributions of state-action pairs $\rho\times\pi$ can be bounded by the error on the samples from $\rho\times\pi$ up to some constant. The single-policy concentrability assumption is strictly weaker than the all-policy concentrability in Assumption 2.6.
2.2 Review of Existing Results
In this subsection, we discuss the direct adaptation of previous results to our setting and demonstrate that their results cannot imply an $\widetilde{O}(\epsilon^{-1})$ sample complexity without all-policy concentrability. Specifically, Xiong et al. (2024) obtained a performance gap upper bound under linear function approximation
$J(\pi^{*})-J(\pi)\leq\big{\|}\mathbb{E}_{\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}_{\bm{\Sigma}_{\text{off}}^{-1}}\eqqcolon\text{RHS},$
where $\bm{\nu}$ is the reference vector, $\bm{\phi}(s,a)\in\mathbb{R}^{d}$ is the feature map, and $\Sigma_{\text{off}}=\sum_{i=1}^{n}\bm{\phi}(s_{i},a_{i})\bm{\phi}(s_{i},a_{i})^{\top}$ is the sample covariance matrix. However, we can show that RHS can be bounded from below by
$\big{\|}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}\sqrt{\lambda_{\min}(\bm{\Sigma}_{\text{off}}^{-1})}$ $=\big{\|}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}\lambda_{\max}(\bm{\Sigma}_{\text{off}})^{-1/2}$
$\geq\big{\|}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}\text{tr}(\bm{\Sigma}_{\text{off}})^{-1/2}$
$=\big{\|}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\bm{\phi}(s,a)]-\bm{\nu}\big{\|}\bigg{(}\sum_{i=1}^{n}\|\bm{\phi}(s_{i},a_{i})\|_{2}^{2}\bigg{)}^{-1/2}$
$=\Omega(n^{-1/2}),$
where $\lambda_{\min}$ and $\lambda_{\max}$ is the minimum and maximum eigenvalue of a matrix, the first inequality holds due to the fact that $\mathbf{x}^{\top}\bm{\Sigma}\mathbf{x}\geq\|\mathbf{x}\|_{2}^{2}\lambda_{\min}(\bm{\Sigma})$ and the second inequality holds due to $\lambda_{\max}(\bm{\Sigma})\leq\text{tr}(\bm{\Sigma})$. Zhao et al. (2024) proposed a two-stage learning algorithm and obtained an $\widetilde{O}(\epsilon^{-1})$ sample complexity for online KL-regularized bandits. The algorithm can be adopted to offline learning by removing the second stage [^3] This can be done by setting the $n$ in their paper to 00. and treat the samples from first stage as the offline dataset. An analogous analysis gives a sample complexity of $\widetilde{O}(D^{2}\epsilon^{-1})$, where $D^{2}$ is the all-policy concentrability.
2.3 Algorithm
In this subsection, we present an offline bandit algorithm, KL-PCB, for KL-regularized contextual bandits in Algorithm 1. KL-PCB first leverages least-square estimator to find a function ${\bar{f}}\in\mathcal{F}$ that minimizes its risk on the offline dataset. In Zhao et al. (2024), such ${\bar{f}}$ is directly applied to construct the estimated policy. In contrast, we construct a pessimistic estimator of $f^{*}$ following the well-known pessimism principle in offline RL (Jin et al., 2021).
Specifically, we consider the bonus term defined as
$\Gamma_{n}(s,a)=\beta D_{\mathcal{F}}\big{(}(s,a),\pi^{\mathsf{ref}}\big{)},\forall(s,a)\in{\mathcal{S}}\times\mathcal{A};$ (2.3)
where
$\beta=\sqrt{128\log\big{(}2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta\big{)}/3n+18\epsilon_{c}}.$ (2.4)
We then obtain our pessimistic estimation ${\widehat{f}}$ by setting ${\widehat{f}}={\bar{f}}-\Gamma_{n}$, which is less than $f^{*}$ with high probability. After obtaining the pessimistic estimation, KL-PCB output the policy $\widehat{\pi}$, which maximizes the estimated objective
$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{f}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
the maximizer of which is the counterpart of 2.2, i.e.,
$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}.$
Algorithm 1 Offline KL-Regularized Pessimistic Contextual Bandit (KL-PCB)
0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{F}$, offline dataset $\mathcal{D}=\{(s_{i},a_{i},r_{i})\}_{i=1}^{n}$
1:  Compute the least square estimation of reward function
${\bar{f}}\in\mathop{\mathrm{argmin}}_{f\in\mathcal{F}}\sum_{(s_{i},a_{i},r_{i})\in\mathcal{D}}\big{(}f(s_{i},a_{i})-r_{i}\big{)}^{2}$
2:  Let ${\widehat{f}}\leftarrow{\bar{f}}-\Gamma_{n}$, where $\Gamma_{n}$ is the bonus term in 2.3
2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}$
2.4 Theoretical Results
The sample complexity for KL-regularized contextual bandits is settled in this subsection. We first give the upper bound of KL-PCB.
Theorem 2.8.
Under Assumption 2.7, for sufficiently small $\epsilon\in(0,1)$, if we set $\Gamma_{n}$ as in (2.3), then $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1})$ suffices to guarantee the output policy $\widehat{\pi}$ of Algorithm 1 to be $\epsilon$-optimal with probability at least $1-\delta$.
Remark 2.9.
Previously, Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 2.6. As a comparison, KL-PCB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 2.7, which is weaker than Assumption 2.6
We also provide the sample complexity lower bound of KL-regularized contextual bandits in the following theorem.
Theorem 2.10.
For any $\epsilon\in(0,1),\eta>0$, and algorithm Alg, there is a KL-regularized contextual bandit instance with $O(1)$ single-policy concentrability such that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
Remark 2.11.
Theorem 2.10 is concluded by plugging $D^{2}\geq D^{2}_{\pi^{*}}$ to Theorem 3.1 of Zhao et al. (2024) and thus the sample complexity lower bound also holds for the KL-regularized contextual bandit instance with $O(1)$ all-policy concentrability.
Remark 2.12.
Theorem 2.10 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual bandits requires at least $\widetilde{\Omega}(\epsilon^{-1})$ samples to output an $\epsilon$-optimal policy. This sample complexity lower bound matches the sample complexity upper bound in Theorem 2.8, indicating that KL-PCB is nearly optimal.
2.5 Proof of Theorem 2.8
In this section, we finish the proof of Theorem 2.8. At a high level, if we consider the regularized objective 1.1 multi-arm bandits, then $P\mapsto{\mathsf{KL}}\left(P\|Q\right)$ is $1$-strongly convex w.r.t. ${\mathsf{TV}}\left(\cdot\|\cdot\right)$ (Polyanskiy and Wu, 2024, Exercise I.37), and thus $J(\pi)$ is strongly concave. Therefore, $J(\pi^{*})-J(\widehat{\pi})$ is possible to be of the order $[{\mathsf{TV}}\left(\pi^{*}\|\widehat{\pi}\right)]^{2}\approx\widetilde{O}(n^{-1})$, pretending that $\pi^{*}$ is the unconstrained maximizer. This intuition guides our analysis for contextual bandits.
We begin the proof with the definition of the event $\mathcal{E}(\delta)$ given $\delta>0$ as
$\mathcal{E}(\delta)\coloneqq\bigg{\{}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Big{[}\big{|}{\bar{f}}-f^{*}\big{|}-\Gamma_{n}\Big{]}(s,a)\leq 0\bigg{\}},$ (2.5)
where $\Gamma_{n}$ is defined in (2.3). Event $\mathcal{E}(\delta)$ holds indicates that the least square estimation ${\bar{f}}$ obtained in Line 1 of Algorithm 1 does not deviate too much from the true function $f^{*}$. More specifically, we have the following lemma, whose proof are deferred to Please refer to Appendix A.1.
Lemma 2.13.
For all $\delta>0$, $\mathcal{E}(\delta)$ holds with probability at least $1-\delta$.
The following covariance-type observation is the first pivot.
Lemma 2.14.
If a bounded random variable $X\leq 0$ almost surely, then $\mathbb{E}[X^{3}]-\mathbb{E}[X^{2}]\mathbb{E}[X]\leq 0$.
Remark 2.15.
While this lemma is elementary, to the best of our knowledge, we are the first to isolate this structure from our non-standard analysis of offline RL, from which the sharp upper bound is derived.
Proof of Lemma 2.14.
We define $Y=-X$. Then it suffices to show that the covariance between $Y$ and $Y^{2}$ is
$\operatorname{\rm Cov}(Y,Y^{2})$ $=\mathbb{E}[Y^{3}]-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
$\geq\big{(}\mathbb{E}[Y^{2}]\big{)}^{3/2}-\mathbb{E}[Y^{2}]\mathbb{E}[Y]$
$=\big{(}\mathbb{E}[Y^{2}]\big{)}\big{(}\sqrt{\mathbb{E}[Y^{2}]}-\mathbb{E}[Y]\big{)}$
$\geq 0,$
where both inequalities follow from Jensen’s inequality. ∎
We further define the following quantities. For all $\gamma\in[0,1]$, we define
$f_{\gamma}\coloneqq\gamma{\widehat{f}}+(1-\gamma)f^{*},$
and based on which we further denote
$\pi_{\gamma}(\cdot|s)\propto\pi^{\mathsf{ref}}(\cdot|s)\exp\big{(}\eta f_{\gamma}(s,\cdot)\big{)},\forall s\in{\mathcal{S}};$
$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\Big{[}\big{(}{\widehat{f}}-{f^{*}}\big{)}^{2}(s,a)\Big{]}.$
The key to our analysis is the monotonicity of the function $G(\gamma)$ in $\gamma$, which is formally stated in the following lemma.
Lemma 2.16.
On event $\mathcal{E}$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
Proof.
For simplicity, we use $\triangle(s,a)$ to denote $\big{(}{\widehat{f}}-{f^{*}})(s,a)$ in this proof. Then we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$ on event $\mathcal{E}$. The most direct way to prove is to take derivative of $G$ with respect to $\gamma$, which corresponds to the policy gradient (Sutton et al., 1999) of $\pi_{\gamma}$ and thus implying a favorable structure. A direct calculation yields that
$G^{\prime}(\gamma)$ $=\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\nabla_{\gamma}\log\pi_{\gamma}(a|s)\triangle(s,a)^{2}\big{]}$
$=\eta\mathbb{E}_{\rho}\mathbb{E}_{a\sim\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{(}\triangle(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\gamma}}[\triangle(s,a^{\prime})]\big{)}\big{]}$
$=\eta\mathbb{E}_{\rho}\Big{[}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{3}(s,a)\big{]}-\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{]}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle(s,a)\big{]}\Big{]}$
$\leq 0,$
where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$, the first equation is derived from standard policy gradient and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.14. ∎
Finally, we need the following lemma to bound the performance difference between two policy w.r.t. the KL-regularized objective $J$. For any given $f:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, we define $\pi_{f}(\cdot|s)\propto\exp\big{(}\eta\cdot f(s,\cdot)\big{)},\forall s\in{\mathcal{S}}$ to facilitate presentation; then $\pi^{*}=\pi_{f^{*}}$ by definition. Please refer to Appendix A.2 for the proof of the lemma.
Lemma 2.17.
Given $f:{\mathcal{S}}\times\mathcal{A}\to\mathbb{R}$, there exist $\gamma\in[0,1]$ such that for $f_{\gamma}=\gamma f+(1-\gamma)f^{*}$ and $\pi_{\gamma}\coloneqq\pi_{f_{\gamma}}$,
$J(\pi^{*})-J(\pi_{f})\leq\eta\mathbb{E}_{(s,a)\sim\rho\times\pi_{\gamma}}\big{[}\big{(}f^{*}-f\big{)}^{2}(s,a)\big{]}.$
Now we are ready to prove Theorem 2.8.
Proof of Theorem 2.8.
Following the proof of Zhao et al. (2024, Theorem 3.3), we know that there exists $\bar{\gamma}\in[0,1]$ such that
$J(\pi^{*})-J(\widehat{\pi})\leq\eta G(\bar{\gamma})\leq\eta G(0),$ (2.6)
where the first inequality holds due to Lemma 2.17 and the second inequality holds due to the event $\mathcal{E}$ and Lemma 2.16. The term $G(0)$ can be further bounded by
$G(0)$ $=\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{f}}-{f^{*}}\big{)}^{2}(s,a)\Big{]}$
$\leq 4\eta\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[\Gamma_{n}^{2}(s,a)]$
$=4\eta\beta^{2}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}),$ (2.7)
where the second inequality holds conditioned on $\mathcal{E}(\delta)$ because of Lemma 2.16, and the last inequality follows from the definition of $\mathcal{E}(\delta)$ together with Line 2. By Lemma 2.13, we know that event $\mathcal{E}$ holds with probability at least $1-\delta$, which finishes the proof. ∎
3 KL-Regularized Contextual Dueling Bandits
In this section, we extend our algorithm to KL-regularized dueling bandit setting, where the learner receives preference feedback instead of absolute signals. We formally introduce setting and extended algorithm. Except the objective, the core setting follows Zhu et al. (2023); Zhan et al. (2023). Our objective follows Xiong et al. (2024); Zhao et al. (2024).
3.1 Problem Setup
We still [^4] We overload some notations in Section 2 by their dueling counterparts for notational simplicity. consider contextual bandits $({\mathcal{S}},\mathcal{A},r,\pi^{\mathsf{ref}})$ where ${\mathcal{S}}$ is the state space, $\mathcal{A}$ is the action space and $r:{\mathcal{S}}\times\mathcal{A}\to[0,1]$ is the reward function. But only relative preference feedback is available, viz., we have an $\mathrm{i.i.d.}$ offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$, where $s_{i}\in{\mathcal{S}}$ is generated from distribution $\rho$ and $a^{1}_{i},a^{2}_{i}\sim\pi^{\mathsf{ref}}$. The binary preference label $y_{i}=1$ indicates $a_{i}^{1}$ is preferred over $a_{i}^{2}$ and 00 for the opposite. Given a state $s$ and two corresponding action $a^{1}$ and $a^{2}$, let $y$ be the preference label such that $y=1$ indicates $a^{1}\succ a^{2}$ and $y=0$ for $a^{2}\succ a^{1}$. In this work we consider the Bradley-Terry Model, where $\mathbb{P}[y=1|s,a^{1},a^{2}]=\sigma(r(s_{i},a^{1}_{i})-r(s_{i},a^{2}_{i}))$, where $\sigma(x)=(1+e^{-x})^{-1}$ is the link function. The objective here identical to (2.1) and thus the optimal policy $\pi^{*}$ also follows the closed form in (2.2). Our goal is still to find an $\epsilon$-optimal policy. To control the complexity of the function class $\mathcal{F}$, we assume that Assumption 2.1 still holds here.
Concentrability
Analogous to Section 2, we need our estimation from offline dataset generalizable to the state-action pairs visited by our obtained policy, and therefore requires similar $D^{2}$-divergence and corresponding policy coverage conditions. However, in dueling bandit, we are unable to observe and estimate the absolute value of rewards. Therefore, intuitively, the best estimation $f$ we can achieve is that for any state $s$ and actions $a^{1},a^{2}$, our estimated $f(s,a^{1})-f(s,a^{2})\approx r(s,a^{1})-r(s,a^{2})$. This implies that there exists some mapping $b:{\mathcal{S}}\to[-1,1]$ such that $f(s,a)-b(s)\approx r(s,a)$ on the offline data. Consequently, it is natural to define the the following modified $D^{2}$-divergence for characterizing how the error $f(s,a)-b(s)-r(s,a)$ generalize to unseen samples.
Definition 3.1.
Given a class of functions $\mathcal{F}\subset({\mathcal{S}}\times\mathcal{A}\to\mathbb{R})$ and some policy $\pi$, let $\mathcal{B}=({\mathcal{S}}\to[-1,1])$ be the function class, define the $D^{2}$-divergence $D^{2}_{\mathcal{F}}((s,a);\pi)$ as
$\sup_{f,g\in\mathcal{F}}\inf_{b\in\mathcal{B}}\frac{\big{(}f(s,a)-g(s,a)-b(s)\big{)}^{2}}{\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi(\cdot|s^{\prime})}[f(s^{\prime},a^{\prime})-g(s^{\prime},a^{\prime})]}.$
A similar definition has been introduced in Zhao et al. (2024, Definition 2.6), which underpins the following two assumptions characterizing the coverage ability of $\pi^{\mathsf{ref}}$.
Equipped with the overloaded $D^{2}$-divergence definition, we are now able to define all-policy concentrability and single concentrability similar as in Section 2.
Assumption 3.2 (All-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D>0$ such that $D^{2}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})$.
Assumption 3.3 (Single-Policy Concentrability).
Given a reference policy $\pi^{\mathsf{ref}}$, there exists $D_{\pi^{*}}>0$ such that $D_{\pi^{*}}^{2}=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}[D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})]$.
Remark 3.4.
Similar single-policy concentrability assumptions have appeared in previous work in offline contextual dueling bandits. Specifically, Huang et al. (2024) defined an $L_{\infty}$-concentrability coefficient $\mathcal{C}_{\infty}^{\pi^{*}}=\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\big{(}\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\big{)}$. Song et al. (2024) introduced a stronger assumption that $\mathcal{C}_{\infty}^{\pi}$ is bounded for all the policy $\pi$ in the KL-ball surrounding $\pi^{*}$. In our Definition 3.1, instead of simply considering the ratio of probability measure, we introduce a more refined eluder-type uncertainty term which also relies on the structure of the model class. As a result, $D_{\pi^{*}}$ can be small even if $L_{\infty}$-concentrability coefficient of $\pi^{*}$ is extremely large as long as the reward model generalizes well after training on the offline dataset. There has also been similar notions of partial coverage for the analysis of model-based RL (Uehara and Sun, 2021; Wang et al., 2024), which fused the information of the transition kernel class to the density-ratio-based formulation.
3.2 Algorithm
We elucidate KL-PCDB for offline KL-regularized contextual dueling bandits, whose pseudocode is summarized in Algorithm 2. KL-PCDB first estimate the ground truth function $f^{*}$ on offline dataset with maximum likelihood estimator (MLE) to estimate a function ${\bar{f}}\in\mathcal{F}$. After that, analogous to Algorithm 1, we adopt the principle of pessimism in the face of uncertainty. Specifically, we define the penalty term
$\Gamma_{n}(s,a)=\beta\sqrt{D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})},$ (3.1)
where
$\beta^{2}=128\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)/3n+18\epsilon_{c}=\widetilde{O}(n^{-1})$ (3.2)
and then subtract it from the MLE ${\bar{f}}$ to obtain a pessimistic estimator ${\widehat{f}}$. KL-PCB then output the policy $\widehat{\pi}$, maximizing the estimated objective
$\widehat{J}(\pi)=\mathbb{E}_{(s,a)\sim\rho\times\pi}\bigg{[}{\widehat{f}}(s,a)-\eta^{-1}\log\frac{\pi(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]},$
the maximizer of which is in closed form as the counterpart of 2.2.
$\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}.$
Algorithm 2 Offline KL-Regularized Pessimistic Contextual Dueling Bandit (KL-PCDB)
0:  regularization $\eta$, reference policy $\pi^{\mathsf{ref}}$, function class $\mathcal{F}$, offline dataset $\mathcal{D}=\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$
1:  Compute the maximum likelihood estimator of the reward function
${\bar{f}}=\mathop{\mathrm{argmin}}_{f\in\mathcal{F}}\sum_{i=1}^{n}\Big{[}y_{i}\log\sigma\Big{(}\big{[}f(s_{i},a^{1}_{i})-f(s_{i},a^{2}_{i})\big{]}\Big{)}+(1-y_{i})\log\sigma\Big{(}\big{[}f(s_{i},a^{2}_{i})-f(s_{i},a^{1}_{i})\big{]}\Big{)}\Big{]}$
2:  Let ${\widehat{f}}(s,a)={\bar{f}}(s,a)-\Gamma_{n}(s,a)$, where $\Gamma_{n}(s,a)$ is the bonus term in 3.1
2:  $\widehat{\pi}(a|s)\propto\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta\cdot{\widehat{f}}(s,a)\big{)}$
3.3 Main Results
In this subsection, we provide our main results for KL-regularized contextual dueling bandits. The sample complexity upper bound of KL-PCDB is presented in the following theorem.
Theorem 3.5.
Under Assumption 3.3, if we set $\Gamma_{n}$ according to Lemma 3.1, then for sufficiently small $\epsilon\in(0,1)$, with probability at least $1-\delta$, $n=\widetilde{O}(\eta D_{\pi^{*}}^{2}\epsilon^{-1})$ is sufficient to guarantee the output policy $\widehat{\pi}$ of Algorithm 2 to be $\epsilon$-optimal.
Remark 3.6.
Zhao et al. (2024) achieved an $\widetilde{O}(\epsilon^{-1})$ sample complexity under Assumption 3.2. Comparing to Zhao et al. (2024), KL-PCDB achieves the same $\widetilde{O}(\epsilon^{-1})$ sample complexity but only requiring Assumption 3.3, which is weaker than Assumption 3.2.
The following theorem provides the sample complexity lower bound for KL-regularized dueling contextual bandits.
Theorem 3.7.
For any $\epsilon\in(0,1)$, $\eta>0$, and offline RL algorithm Alg, there is a KL-regularized contextual dueling bandit instance with $O(1)$ single-policy concentrability such that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ samples to return an $\epsilon$-optimal policy.
Proof.
See Appendix A.3. ∎
Remark 3.8.
Theorem 3.7 can be concluded by plugging $D^{2}\geq D^{2}_{\pi^{*}}$ into Zhao et al. (2024, Theorem 4.3). This implies that the sample complexity lower bound in Theorem 3.7 also holds for KL-regularized dueling contextual bandit with all-policy concentrability.
Remark 3.9.
Theorem 3.7 shows that when $\epsilon$ is sufficiently small, any algorithm for offline KL-regularized contextual dueling bandit requires at least $\widetilde{\Omega}(\epsilon^{-1})$ samples to output a $\epsilon$-optimal policy. This sample complexity lower bound matches the sample complexity upper bound in Theorem 3.5, indicating that KL-PCDB is nearly optimal.
3.4 Proof of Theorem 3.5
The proof follows the proof in Section 2. At the beginning, we first define the event $\mathcal{E}(\delta)$ given $\delta>0$ as
$\mathcal{E}(\delta)\coloneqq\Big{\{}\exists\ b:{\mathcal{S}}\to[-1,1],\forall(s,a)\in{\mathcal{S}}\times\mathcal{A},\big{|}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{|}\leq\Gamma_{n}(s,a)\Big{\}}.$ (3.3)
Here, $\Gamma_{n}$ is defined in (3.1). We abuse the notation and define $b(\cdot)$ as
$b=\mathop{\mathrm{argmin}}_{\mathcal{B}}\sup_{(s,a)\in{\mathcal{S}}\times\mathcal{A}}\Phi_{b}(s,a)-\Gamma_{n}(s,a),$ (3.4)
where $\Phi_{b}(s,a)=\big{|}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{|}$ and when $\mathcal{E}$ holds, for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we have $\Phi_{b}(s,a)\leq\Gamma_{n}(s,a)$ This indicates that the least square estimation ${\bar{f}}$ obtained in Line 1 of Algorithm 2, after adjusted by some bias function $b$, is close to the true function $f^{*}$. The following lemma shows that this event holds with high probability. Please refer to Appendix B.1 for the proof.
Lemma 3.10.
For any $\delta>0$, $\mathbb{P}(\mathcal{E}(\delta))\geq 1-\delta$.
We overload the following quantities. For any $\gamma\in[0,1]$ and $(s,a)\in{\mathcal{S}}\times\mathcal{A}$, we define
$f_{\gamma}(s,a)\coloneqq\gamma({\widehat{f}}(s,a)-b(s))+(1-\gamma)f^{*}(s,a).$
Furthermore, we introduce the following quantities
$\pi_{\gamma}(\cdot|\cdot)=\pi_{f_{\gamma}}(\cdot|\cdot)\propto\pi^{\mathsf{ref}}(\cdot|\cdot)\exp\big{(}\eta f_{\gamma}(\cdot,\cdot)\big{)},$
$G(\gamma)\coloneqq\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\big{]},$
where $b(\cdot)$ is defined in (3.4). We still have the monotonicity of the function $G(\gamma)$, which is characterized by the following lemma.
Lemma 3.11.
On event $\mathcal{E}(\delta)$, $0\in\mathop{\mathrm{argmax}}_{\gamma\in[0,1]}G(\gamma)$.
Proof.
For simplicity, we use $\triangle(s,a)$ to denote ${\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)$ in this proof. Then on event $\mathcal{E}(\delta)$, we know that $\triangle(s,a)\leq 0$ for all $(s,a)\in{\mathcal{S}}\times\mathcal{A}$. Taking derivatives of $G$ w.r.t., $\gamma$ directly, we conclude that for all $\gamma\in[0,1]$,
$G^{\prime}(\gamma)$ $=\eta\mathbb{E}_{\rho}\mathbb{E}_{a\sim\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{(}\triangle(s,a)-\mathbb{E}_{a^{\prime}\sim\pi_{\gamma}}[\triangle(s,a^{\prime})]\big{)}\big{]}$
$=\eta\mathbb{E}_{\rho}\Big{[}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{3}(s,a)\big{]}-\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle^{2}(s,a)\big{]}\mathbb{E}_{\pi_{\gamma}}\big{[}\triangle(s,a)\big{]}\Big{]}$
$\leq 0,$
where $\mathbb{E}_{\rho}$ is the shorthand of $\mathbb{E}_{s\sim\rho}$, $\mathbb{E}_{\pi_{\gamma}}$ is the shorthand of $\mathbb{E}_{a\sim\pi_{\gamma}}$ and the inequality holds conditioned on the event $\mathcal{E}(\delta)$ due to Lemma 2.14. ∎
Finally, we have the proposition that adding some bias term $b:{\mathcal{S}}\to\mathbb{R}$ does not affect the resulting policy.
Proposition 3.12.
Let $b:{\mathcal{S}}\to\mathbb{R}$ be some bias function, then for all $f\in\mathcal{F}$ we have $J(\pi_{f})=J(\pi_{f-b})$, where $(f-b)(s,a)=f(s,a)-b(s)$.
Please see Appendix B.2 for the proof of Proposition 3.12. Now we are ready to prove Theorem 3.5.
Proof of Theorem 3.5.
We proceed the proof under the event $\mathcal{E}(\delta)$. By Proposition 3.12, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{f}}})$
$=J(\pi^{*})-J(\pi_{{\widehat{f}}-b}).$
Consequently, there exist some $\gamma\in[0,1]$ and $b:{\mathcal{S}}\to[-1,1]$ such that
$J(\pi^{*})-J(\widehat{\pi})$ $=J(\pi^{*})-J(\pi_{{\widehat{f}}-b})$
$\leq\eta\mathbb{E}_{\rho\times\pi_{\gamma}}\big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\big{]}$
$=\eta G(\gamma),$ (3.5)
where the inequality holds due to Lemma 2.17. Under event $\mathcal{E}(\delta)$, we know that ${\widehat{f}}(s,a)-b(s)\leq{f^{*}}(s,a)$. Together with Lemma 3.11, we obtain $G(\gamma)\leq G(0)$. Therefore, we know that
$J(\pi^{*})-J(\widehat{\pi})$ $\leq G(0)$ (3.6)
$=\eta\mathbb{E}_{\rho\times\pi^{*}}\Big{[}\big{(}{\widehat{f}}(s,a)-b(s)-{f^{*}}(s,a)\big{)}^{2}\Big{]}$
$\leq 4\eta\mathbb{E}_{\rho\times\pi^{*}}\Gamma_{n}^{2}(s,a)$
$=4\eta\beta^{2}\mathbb{E}_{\rho\times\pi^{*}}\big{[}D^{2}_{\mathcal{F}}((s,a);\pi^{\mathsf{ref}})\big{]}$
$=\widetilde{O}(\eta D_{\pi^{*}}^{2}n^{-1}),$ (3.7)
where the inequality holds due to the definition of $\mathcal{E}(\delta)$. Plugging (3.7) into (3.5), we know that $J(\pi^{*})-J(\widehat{\pi})$ has upper bound $\widetilde{O}(D_{\pi^{*}}^{2}n^{-1})$. By Lemma 3.10, event $\mathcal{E}$ with probability at least $1-\delta$, which concludes the proof. ∎
4 Conclusions and Future Work
In this work, we settled the sample complexity of KL-regularized offline contextual bandits. We first propose KL-PCB, which follows the principle of pessimism for reward estimation and yields the optimal policy accordingly. Our analysis leverages the curvature of KL-regularized objectives and integrates pessimism with a newly identified covariance-based observation, which in turn enables a neat refinement of a mean-value-type argument the its extreme point, showing that KL-PCB achieves a nearly optimal sample complexity under single-policy concentrability. The pair of novel techniques are decoupled from tricky algorithmic tweaks, and thus might be of independent interest. We then extended KL-PCB to offline KL-regularized contextual dueling bandit problem and introduced KL-PCDB, the analysis of which manifests the near-optimality of KL-PCDB under single-policy concentration. Currently, our focus is solely on the reverse-KL-regularized objective. It would be an interesting future direction to study more general regularization techniques like $f$-divergence (Csiszár, 1967) regularization, which includes both forward and reverse KL regularizations.
Appendix A Missing Proofs from Section 2
A.1 Proof of Lemma 2.13
We first provide the following lemmas.
Lemma A.1 (Zhao et al. 2024, Lemma C.1).
For any policy $\pi$ and state-action pairs $\{(s_{i},a_{i})\}_{i=1}^{m}$ generated i.i.d. from $\rho\times\pi$, and $\epsilon_{c}<1$, with probability at least $1-\delta$, for any $f_{1}$ and $f_{2}$ we have
$\mathbb{E}_{\rho\times\pi}\big{[}\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}\big{]}\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}f_{1}(s_{i},a_{i})-f_{2}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c},$
where $\mathcal{N}_{\mathcal{F}}(\epsilon_{c})$ is the $\epsilon_{c}$-covering number of $\mathcal{F}$.
Lemma A.2 (Zhao et al. 2024, Lemma C.2).
For arbitrary policy $\pi$ and dataset $\{(s_{i},a_{i},r_{i})\}_{i=1}^{m}$ generated i.i.d., from the product of $\pi$, $\rho$ and the Bradley-Terry Model; let ${\bar{f}}$ be the least square estimator of $f^{*}$, then for any $10<\epsilon_{c}<$ 1 and $\delta>0$, with probability at least $1-\delta$ we have
$\sum_{i=1}^{n}\big{(}{\bar{f}}(s_{i},a_{i})-f^{*}(s_{i},a_{i})\big{)}^{2}\leq 16\log(a\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+4n\epsilon_{c}.$
Proof of Lemma 2.13.
We have the following inequality
$\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}$ $=\frac{\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$
$\leq\sup_{f_{1},f_{2}\in\mathcal{F}}\frac{\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}}{\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}f_{1}(s,a)-f_{2}(s,a)\big{)}^{2}\big{]}}\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$
$=D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]},$ (A.1)
where the inequality holds by taking supremum to $f_{1},f_{2}\in\mathcal{F}$. Now we have
$\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)\big{)}^{2}\big{]}$ $\leq\frac{2}{n}\sum_{i=1}^{n}\big{(}{\bar{f}}(s_{i},a_{i})-f^{*}(s_{i},a_{i})\big{)}^{2}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
$\leq\frac{2}{n}\big{[}16\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+4n\epsilon_{c}\big{]}+\frac{32}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+10\epsilon_{c}$
$=\frac{128}{3n}\log(2\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+18\epsilon_{c},$ (A.2)
where the first inequality holds due to Lemma A.1 and second holds due to Lemma A.2. Plugging (A.1) into (A.1) and setting $\epsilon_{c}=O(n^{-1})$ complete the proof. ∎
A.2 Proof of Lemma 2.17
This proof is extracted from the proof of Zhao et al. (2024, Theorem 3.3) and we present it here for completeness. By definition of our objective in (2.1), we have
$J(\pi^{*})-J(\pi_{f})$
$\quad=\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}f^{*}(s,a)-\eta^{-1}\log\frac{\pi^{*}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}-\mathbb{E}_{(s,a)\sim\rho\times\pi_{f}}\bigg{[}f^{*}(s,a)-\frac{1}{\eta}\log\frac{\pi_{f}(a|s)}{\pi^{\mathsf{ref}}(a|s)}\bigg{]}$
$\quad=\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi^{*}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta f^{*}(s,a)\bigr{)}}{\pi^{*}(a|s)}\bigg{]}-\frac{1}{\eta}\mathbb{E}_{(s,a)\sim\rho\times\pi_{f}}\bigg{[}\log\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta f^{*}(s,a)\bigr{)}}{\pi_{f}(a|s)}\bigg{]}$
$\quad=\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{f^{*}}(s)\big{]}-\frac{1}{\eta}\mathbb{E}_{s\sim\rho}\big{[}\log Z_{f}(s)\big{]}-\mathbb{E}_{s\sim\rho}\bigg{[}\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\big{(}f^{*}(s,a)-f(s,a)\big{)}\bigg{]},$
where for all $f\in\mathcal{F}$ we define $Z_{f}(\cdot)$ as follows,
$Z_{f}(\cdot)\coloneqq\sum_{a\in\mathcal{A}}\pi^{\mathsf{ref}}(a|\cdot)\exp\big{(}\eta f(\cdot,a)\big{)}.$
We further denote $\Delta(s,a)=f(s,a)-f^{*}(s,a)$ and $H_{s}(f)=\log Z_{f}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\Delta(s,a)$. It worth noticing that $\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}(f^{*})-H_{s}(f)]=J(\pi^{*})-J(\pi_{f})$. Now we take derivative of $H$ with respect to $\Delta(s,a)$,
$\frac{\partial H_{s}(f)}{\partial\Delta(s,a)}$ $=\frac{\partial}{\partial\Delta(s,a)}\bigg{[}\log Z_{f}(s)-\eta\sum_{a\in\mathcal{A}}\pi_{f}(a|s)\cdot\Delta(s,a)\bigg{]}$
$=\frac{1}{Z_{f}(s)}\cdot\pi^{\mathsf{ref}}(a|s)\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}\cdot\eta-\eta\cdot\pi_{f}(a|s)$
$\quad-\eta^{2}\cdot\Delta(s,a)\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}}{Z_{f}(s)}+\eta^{2}\cdot\Delta(s,a)\cdot\frac{\bigl{[}\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}\bigr{]}^{2}}{[Z_{f}(s)]^{2}}$
$\quad+\eta\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\frac{\pi^{\mathsf{ref}}(a^{\prime}|x)\cdot\exp\bigl{(}\eta\cdot f(s,a^{\prime})\bigr{)}}{Z_{f}(s)}\cdot\eta\cdot\Delta(s,a^{\prime})\cdot\frac{\pi^{\mathsf{ref}}(a|s)\cdot\exp\bigl{(}\eta\cdot f(s,a)\bigr{)}}{Z_{f}(s)}$
$=-\eta^{2}\pi_{f}(a|s)\Delta(s,a)+\eta^{2}[\pi_{f}(a|s)]^{2}\cdot\Delta(s,a)+\eta^{2}\sum_{a^{\prime}\in\mathcal{A}\backslash\{a\}}\pi_{f}(a^{\prime}|x)\pi_{f}(a|s)\Delta(s,a^{\prime}).$
Therefore, by mean value theorem, there exists $\gamma\in[0,1]$ and $f_{\gamma}=\gamma f+(1-\gamma)f^{*}$ such that
$H_{s}(f)-H_{s}(f^{*})$ $=-\eta^{2}\gamma\sum_{a\in\mathcal{A}}\pi_{f_{\gamma}}(a|s)\Delta(s,a)^{2}+\gamma\eta^{2}\sum_{a_{1}\in\mathcal{A}}\sum_{a_{2}\in\mathcal{A}}\pi_{f_{\gamma}}(a_{1}|x)\pi_{f_{\gamma}}(a_{2}|x)\Delta(s,a_{1})\Delta(s,a_{2})$
$=-\eta^{2}\gamma\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]}+\gamma\eta^{2}\Big{(}\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}\big{]}\Big{)}^{2}$
$\geq-\eta^{2}\mathbb{E}_{a\sim\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]},$
where the inequality holds by omitting the second term and $\gamma\leq 1$. Now taking expectation over $\rho$, we have
$J(\pi^{*})-J(\pi_{f})$ $=\eta^{-1}\mathbb{E}_{s\sim\rho}[H_{s}(f^{*})-H_{s}(f)]$
$\leq\eta\mathbb{E}_{(s,a)\sim\rho\pi_{f_{\gamma}}}\big{[}\big{(}f^{*}(s,a)-f(s,a)\big{)}^{2}\big{]},$
which concludes the proof.
A.3 Proof of Theorem 2.10
Proof of Theorem 2.10.
We consider the bandit instance $\mathcal{K}$ and corresponding function class $\mathcal{F}$ constructed in the proof of Zhao et al. (2024, Theorem 3.1). Since the constructed $\pi^{\mathsf{ref}}(\cdot|s)$ in $\mathcal{K}$ is a uniform distribution for any $s\in{\mathcal{S}}$, the all-policy concentrability coefficient $D^{2}$ of $\mathcal{K}$ defined in Assumption 2.6 satisfies $D^{2}=O(1)$, which indicates that the single-policy concentrability coefficient $D_{\pi^{*}}^{2}$ defined in Assumption 2.7 satisfies $D_{\pi^{*}}^{2}\leq D^{2}=O(1)$. Thus, the proof of Zhao et al. (2024, Theorem 3.1) implies that any Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ to obtain an $\epsilon$-optimal policy. ∎
Appendix B Missing Proofs from Section 3
B.1 Proof of Lemma 3.10
The following on-poicy concentration bound is the prerequisite.
Lemma B.1 (Zhao et al. 2024, Lemma D.4).
Consider a offline dataset $\{(s_{i},a^{1}_{i},a^{2}_{i},y_{i})\}_{i=1}^{n}$ generated from the product of the context distribution $\rho\in\Delta({\mathcal{S}})$, policy $\pi\in\Delta(\mathcal{A}|{\mathcal{S}})$, and the Bradley-Terry Model defined in Section 3.1. Suppose ${\bar{f}}$ is the result of MLE estimation of Algorithm 2, and we further define $b(s)=\mathbb{E}_{a\sim\pi(\cdot|s)}\big{[}{\bar{f}}(s,a)-f^{*}(s,a)\big{]}$, then with probability at least $1-2\delta$, we have
$\mathbb{E}_{s,a\sim\rho\times\pi}\big{[}\big{(}{\bar{f}}(s,a)-f^{*}(s,a)-b(s)\big{)}^{2}\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$
Proof of Lemma 3.10.
From Lemma B.1, we have that with probability at least $1-\delta$, it holds that
$\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}\leq O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)}.$ (B.1)
It further holds true that for some $b:{\mathcal{S}}\to\mathbb{R}$
$D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}}))\cdot\mathbb{E}_{s\sim\rho}\operatorname{{\rm Var}}_{a\sim\pi^{\mathsf{ref}}(\cdot|s)}\big{[}{\bar{f}}(s,a)-f^{*}(s,a)\big{]}\geq\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}.$ (B.2)
Substituting (B.1) into (B.2), we have
$\inf_{b}\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}$
$\quad=\inf_{b}\frac{\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}}{\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}}\mathbb{E}_{s^{\prime}\sim\rho}\operatorname{{\rm Var}}_{a^{\prime}\sim\pi^{\mathsf{ref}}(\cdot|s^{\prime})}\big{[}{\bar{f}}(s^{\prime},a^{\prime})-f^{*}(s^{\prime},a^{\prime})\big{]}$
$\quad\leq D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})\mathbb{E}_{\pi^{\mathsf{ref}}}\big{[}\big{(}{\bar{f}}(s,a)-b(s)-f^{*}(s,a)\big{)}^{2}\big{]}$
$\quad\leq D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})O\bigg{(}\frac{1}{n}\log(\mathcal{N}_{\mathcal{F}}(\epsilon_{c})/\delta)+\epsilon_{c}\bigg{)},$
where the first inequality holds due to the definition of $D_{\mathcal{F}}^{2}((s,a),\pi^{\mathsf{ref}})$ and the last inequality holds due to Lemma B.1. ∎
B.2 Proof of Proposition 3.12
Proof of Proposition 3.12.
For any fixed state $s\in{\mathcal{S}}$, we have for any $a\in\mathcal{A}$ that,
$\pi_{f}(a|s)$ $=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta f(s,a)\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta f(s,a^{\prime})\big{)}}$
[…]
$=\frac{\pi^{\mathsf{ref}}(a|s)\exp\big{(}\eta[f(s,a)-b(s)]\big{)}}{\sum_{a^{\prime}\in\mathcal{A}}\pi^{\mathsf{ref}}(a^{\prime}|s)\exp\big{(}\eta[f(s,a^{\prime})-b(s)]\big{)}}$
$=\pi_{f-b}(a|s),$
which indicates that $\pi_{f}=\pi_{f-b}$. This immediately leads to $J(\pi_{f})=J(\pi_{f-b})$. ∎
B.3 Proof of Theorem 3.7
Proof.
Notice that Assumptions 3.2 and 3.3 imply $D^{2}\geq D^{2}_{\pi^{*}}$, which enables us to adapt Zhao et al. (2024, Theorem 4.3) to our scenario. In particular, we consider the bandit instance $\mathcal{K}$ and the corresponding function class $\mathcal{F}$ constructed in Zhao et al. (2024, Theorem 4.3). By Theorem 4.3 of Zhao et al. (2024), we know that the all-policy concentrability coefficient $D^{2}$ of $\mathcal{K}$ satisify $D^{2}=O(1)$, which indicates that its single-policy concentrability coefficient $D_{\pi^{*}}=O(1)$. Theorem 4.3 of Zhao et al. (2024) further indicates that Alg requires at least $\Omega\big{(}\min\{{\eta\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon},{\log\mathcal{N}_{\mathcal{F}}(\epsilon)}/{\epsilon^{2}}\}\big{)}$ to obtain an $\epsilon$-optimal policy, which finishes the proof. ∎